OpenStudy (anonymous):

solve x + 7 = -2(4x + 10)

OpenStudy (anonymous):

x + 7 = -2(4x+10) x + 7 = -8x - 20 +27 = -9x - Subtract negative 20 from right side and subtract 'X' from left side +27 / -9 = -3 - Divide -9 from both sides to remove -9 from 'X' to solve for 'X' x = -3